### v1

# Inga Strümke5, Vince I. Madai6,7, Tobias Budig8, Ali Sunyaev9, and Adam Hilbert10 On behalf of the VALIDATE consortium

Abstract
The joint implementation of Federated learning (FL) and Explainable artificial intelligence (XAI) will allow training models from distributed data and explaining their inner workings while preserving important aspects of privacy. Towards establishing the benefits and tensions associated with their interplay, this scoping review maps those publications that jointly deal with FL and XAI, focusing on publications where an interplay between FL and model interpretability or post-hoc explanations was found.
In total, 37 studies met our criteria, with more papers focusing on explanation methods (mainly feature relevance) than on interpretability (mainly algorithmic transparency). Most works used simulated horizontal FL setups involving 10 or fewer data centers. Only one study explicitly and quantitatively analyzed the influence of FL on model explanations, revealing a significant research gap. Aggregation of interpretability metrics across FL nodes created generalized global insights at the expense of node-specific patterns being diluted. 8 papers addressed the benefits of incorporating explanation methods as a component of the FL algorithm.
Studies using established FL libraries or following reporting guidelines are a minority. More quantitative research and structured, transparent practices are needed to fully understand their mutual impact and under which conditions it happens.
Index Terms:
Machine Learning, Artificial Intelligence (AI), Explainable AI, Model Interpretability, Federated Learning, Data Privacy Preservation.
I Introduction
Figure 1: Schematic overview of a main research question: do explanations differ between federated models and centralized models?
The development of trustworthy AI systems depends on multiple different ethical principles like privacy preservation and explicability [1, 2]. Ensuring that AI systems address such principles is crucial for maintaining trust, compliance, and ethical standards. Such principles are critical in data-sensitive applications such as banking [3] and healthcare [4, 5].
Privacy concerns are the main reason why healthcare data for AI models, such as patient outcome prognosis, mainly originate from single institutions. Anonymizing extensive healthcare data effectively is challenging due to the high risk of re-identifying patients [6]. Allowing models to be trained across multiple institutions without sharing raw data, for example, with federated learning (FL), addresses this issue. FL has gained prominence as a solution to train models from distributed data while preserving these privacy aspects [7, 8]. Specifically, FL enables the training of machine learning (ML) models from distributed data that are available to different users or institutions, without data being explicitly transferred. Therefore, FL solves certain privacy-related governance issues as it restricts access to sensitive information from each institution while still gaining insights from all available data [9, 10]. Using data from different institutions can lead to more diverse training data, which reduces the bias of ML models and improves generalizability compared to models trained on homogeneous data [11]. The three main categories within FL are horizontal FL (HFL), vertical FL (VFL), and federated transfer learning (FTL). HFL involves data scattered across the sample space, where multiple parties have different subsets of data points but all data have the same feature space. VFL allows learning from data distributed across different different feature spaces for the same data sample. FTL focuses on transferring knowledge across different federated learning settings [12].
Another ethical principle often advocated by ethicists and policymakers in high-risk AI domains such as healthcare is explicability. This means that: a) processes must be transparent, b) the capabilities and purpose of AI systems must be openly communicated, and c) one must be able to explain the decisions and predictions made by AI systems to those directly or indirectly affected by them [1, 2]. When a model’s inference or decision-making process is transparent and understandable, it is more likely to be trusted and accepted by stakeholders, such as customers, regulators, and users [13].
As stated by the European Union’s regulations, aspects (a) and (c) of explicability, whose technical manifestation is known as explainability [14], need to be provided through additional information — next to the performance of the AI system — on how an AI system arrives at a certain output [15]. The concept of explainable artificial intelligence (XAI) aims to explain the inner workings of ML algorithms and increasing model comprehension. XAI is thereby concerned with the development of algorithms and methods that achieve explainability [16]. Explicability can be achieved by either training interpretable ML models, or by using post-hoc explanation methods (i.e., (a) and (c) respectively regarding the European Union guideline).
Unfortunately, XAI literature has not agreed on one definition of explainability. The terms “explainability”, “explicability”, and “interpretability” are often used interchangeably [17]. In this work, we adopt the definitions provided by [15], which include interpretable modeling and explanation methods, as detailed in the following paragraphs.
An ML model is said to be interpretable if its decision process can be inherently and intuitively understood by the intended user [15]. The word ‘inherently‘ is crucial in this definition, as interpretability of a given model is defined as a passive property that reflects the level at which such a model makes sense for a human observer [18] without the need to perform extra computation. Examples of interpretable ML models are linear and logistic regression models, where variable importance can be inferred by investigating model weights, or decision trees, which can be intuitively followed by humans [15].
In contrast, post-hoc explanation methods try to explain non-interpretable models actively by performing additional computations with the intent of clarifying or detailing their internal functions [18, 15]. Such computations include evaluating or calculating gradients of the original model for a set of differently perturbed inputs [19]. With that, AI systems can calculate the importance of features [20], the dependence on feature variations of the output [21], or the minimal input change that causes a label flip [22]. An explanation method, therefore, is a comprehensible interface between humans and a decision maker while simultaneously providing approximations of the decision-maker [23].
Extensive research has been conducted on FL and XAI separately. Despite the importance of their combined use for creating trustworthy AI systems, it remains unclear whether these two technologies complement each other, present conflicting requirements, or can be addressed separately. FL and XAI coexist in many works, but few works discuss the impact of one on the other. Furthermore, the existing literature inadequately addresses the methodologies, challenges, and outcomes related to the integration of FL and XAI, leaving a critical gap in understanding their mutual influence.
On one hand, XAI can mitigate the risks associated with training ML systems in federated settings. Learning from heterogeneous data can inadvertently lead to spurious correlations learned by the ML model, as the decentralized data might not represent the overall distribution and contain biases specific to subsets of the data. Explainability can help practitioners determine whether the model’s predictions are based on valid patterns rather than misleading correlations, not propagating undesired biases and, therefore, adhering to ethical standards. Additionally, explainability can help identify malicious agents trying to poison the FL process.
On the other hand, the potential risks associated with the interplay between these technologies must be documented. There is no unified effort to determine, for example, whether FL leads to reduced interpretability or less accurate explanations. Also, explanations could make the FL network more vulnerable to attacks. Moreover, preliminary research [24] indicates differences in model explanations between FL and traditional centralized models, suggesting unique challenges in applying XAI in FL contexts. This underscores the need for further research on how FL influences the contribution of input variables and the effectiveness of XAI methods.
To understand the interplay between FL and XAI, it is crucial to consider the findings, setups, and conditions under which these studies have been conducted. This scoping review aims to map both experimental and methodological studies that examine the interaction between FL and XAI. Since FL can potentially influence both model interpretability and post-hoc explanations, this review paper investigates both concepts of XAI. It focuses on research articles that either propose FL methods incorporating explainability, discuss practical applications where FL and XAI are jointly used, or analyze the interplay between these technologies through empirical studies. Given the fragmented and emerging nature of research at this intersection, this review provides a comprehensive overview of current work, identifies key concepts and evidence types, and highlights gaps to guide future research efforts.
The remainder of this paper is structured as follows. Sec. II introduces the related work. Sec. III describes the methodology we followed. A more detailed analysis of the interplay between FL and XAI found in the reviewed papers is given in Sec. IV. A quantitative analysis of the extracted information is shown in Sec. V. The results are discussed in Sec. VI, and Sec. VII concludes the paper. The Appendix contains summarized information about how all included papers jointly with FL and XAI.
II Related Work
Review papers on the fields of XAI and FL separately are very common, and there are even meta-reviews [25] and systematic reviews from different points of view (see, e.g. [26, 27] regarding FL applications for biomedical data). An extensive review on XAI methods can be found in [16] or [28]. Even though, there are many fewer reviews concerning the joint application of XAI and FL, none analyzes the interplay in depth, which motivates our work.
II-A Reviews about FL and XAI
The review in [29] introduces Fed-XAI, which involves learning interpretable models in federation, in addition to explanation methods applied on any federated trained model. They used [18] as a source for XAI taxonomy and highlight a general diversity in definitions used in the studied literature. Additionally, they gave an overview of the “current status in Fed-XAI” by discussing a set of works relating to the two concepts with no analysis on the interplay between FL and XAI.
The review in [30] gathers relevant works concerning the “FED-XAI” concept, defined by them as a discipline “which aims to bring together these two approaches into one”. Differently from our search terms, which include variations of the words “explanation” and “interpretability”, their search is limited to the explicit string “FED-XAI”, returning a reduced number of publications. The authors do not provide a comparative analysis of the interplay between FL and XAI among the found papers.
The paper [31] surveys “interpretable federated learning” and propose an interpretable FL taxonomy that enables learning models to explain prediction results, support model debugging, and provide insights into data owner contributions. The survey analyzes representative interpretable FL approaches, commonly adopted performance evaluation metrics, and future research directions. The main focus is on how to use interpretable methods to improve the FL process, but not on the impact of FL on the explanations of the predictions.
In contrast to these works, our review 1) applies a scoping review methodology, 2) provides a more detailed set of results, 3) summarizes how the reviewed works jointly deal with FL and XAI, and 4) analyzes the findings on the interplay between them.
II-B Contribution-aware FL
A small but salient subset of the approaches investigating the intersection of FL and XAI has utilized feature relevance methods (e.g. Shapley values) to measure the contribution of each client participating in a FL process. These approaches, commonly referred to as contribution-aware FL, aim to incentivize participants to engage during model training and to fairly distribute rewards based on contributions. One of the first approaches to contribution-aware FL suggested using Shapley values to interpret contributions in FL networks in [32]. The contributions are diverse, ranging from node liability [33] to biases within data sets [34]. Recent approaches have also extended SHAP-based [20] contribution determination to provide visualizations to evaluate data privacy [35] and to improve the reliability of prediction via clustering patients [36]. Such methods do not necessarily constitute an influence of XAI on the FL process and are therefore not central to the present review.
III Methods
The present scoping review addresses the concept of joint application of FL and XAI and their potential interplay. Two research types have been addressed, namely the methodological advances and the experimental results.
We aimed at examining the range and nature of research activity on the topic, summarizing the research findings, and identifying research gaps, following a standard scoping review methodology [37]. The presence of two subconcepts of explainability (interpretable models and post-hoc explanation), the two research types (methodological and experimental), and the diversity of joint approaches to FL and XAI increased the complexity of the study. Understanding and summarizing the diverse approaches required a deeper analysis of each included paper. A more detailed description of the methodology is available in the pre-published research protocol [38].
In our survey, the time frame of the published papers was between 2019 and April 2023 (date of protocol pre-publication). No relevant papers were found before 2019.
III-A Research questions
The research questions of this work address both the training of interpretable ML models using FL, and the explanation methods applied to ML models trained via FL. Furthermore, the questions were categorized into two groups:
III-A1 Methodological advances
  • Which interpretable ML models can be trained via FL?
  • What are existing methods to train interpretable ML models via FL?
  • Are there explanation methods that take into account that the ML model was trained via FL?
  • Has any work proposed an FL method that takes into account the ulterior application of explanation methods?
  • For what subtypes of FL have methods been proposed to a) learn interpretable ML models; b) explain the outputs of ML models?
III-A2 Experimental results
  • In which contexts (fields of application) have results been obtained regarding FL and XAI?
  • Under what conditions does training a model via FL affect a) its interpretability? b) the obtained explanations?
  • Has any work compared the effects of FL on the interpretability of the resulting models against a centralized learning algorithm?
  • Has any work compared the explanations obtained from a model trained via FL against a centralized learning algorithm?
We used the following search engines: Google Scholar, IEEExplore, PubMed, Scopus, and Web of Science. The terminology issue described in the introduction influenced the search strategy, which aimed at identifying all papers mentioning FL jointly with a word related to either interpretability or explainability. The search terms are detailed in Table I.
Library/Engine Search string
IEEEXplore (“federated”) AND (learning) AND ((explainable) OR (interpretable) OR (explaining) OR (explainability) OR (interpreting) OR (interpretability) OR (interpret))
Google Scholar allintitle: (explainable OR explainability OR explaining OR interpretability OR interpretable OR interpret OR intepreting) federated learning
PubMED (“federated”) AND (learning) AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)
Scopus TITLE-ABS-KEY(“federated” AND learning AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)) AND (LIMIT-TO(DOCTYPE, “ar”) OR LIMIT-TO(DOCTYPE, “cp”))
WebOfScience Federated Learning AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)
TABLE I: Search terms
III-B Inclusion/exclusion criteria
We initially identified all research articles published in peer-reviewed conferences and journals and pre-published works made available in preprint services such as arXiv over the past three years. We required that the full text was available.
To determine their inclusion in the review, papers had to either discuss in depth the relation and interaction between XAI and FL, or apply both technologies jointly in a practical setting. To rule out papers that mentioned both technologies without detailed discussion or practical application, we performed a screening procedure, in which the article must answer positively to at least one of the following screening questions:
  • SQ1) Does the paper propose a novel method integrating FL and XAI?
  • SQ2) Does the paper report results from experiments applying FL and XAI in a real dataset?
  • SQ3) Does the paper discuss or assess the impact of FL on explanations or model interpretability?
The exclusion criteria were defined as follows. Theses were excluded. Reviews/meta-reviews/surveys were excluded. Preprints published more than 3 years before the protocol prepublication, for which a final published version did not exist, were excluded.
III-C Extraction
To facilitate efficient and parallel data extraction, we utilized a shared online spreadsheet. All authors were involved in the data extraction which allowed us double-coding of all information. The extracted items were detailed in the protocol [38], and a summarized listed is provided here:
  • Study identification: title, authors, publication year, publication outlet.
  • Nomenclature: whether the paper provides or cites a definition of explainability, and whether the nomenclature regarding explainability coincides with the one introduced here.
  • Nature of the paper: whether it is general/theoretical or applied; if applied, what field it is applied on, and whether the proposed idea is practically validated in the field of application.
  • Data characteristics: information modality, type and amount (number of unique data samples) of the data used in the experiments.
  • FL-specific characteristics: type of FL (HFL, VFL, FTL) used, setup (simulated or not, number of data centers), and which FL library is used (e.g., FLWR, openFL).
  • XAI-specific characteristics: whether an interpretable model or an explanation method is used [^1] If the nomenclature in a papers does not agree with the one mentioned in the Introduction, we identify whether it uses explanation methods or interpretable models according to our understanding., its type, and the specific explanation method (e.g.: SHAP, GradCAM [39]) or way of interpreting the model (e.g., weights at the first layer of DNN).
  • Interplay between FL and XAI: a paragraph summarizing how the paper deals jointly with FL and XAI; influence of FL on explanations or interpretability (e.g. significant differences in variable importance ratings between centrally learned and model trained via FL); influence of XAI onto federated training (e.g., modified merging step in central server after collecting locally trained models); and whether this influence is quantified, and how.
  • Methodology notes: how the novel method is designed (e.g., optimization-based approach), and how rigorous the methodology is.
During both the screening and the extraction process, double coding was practiced. Each screened paper was independently assessed by two different authors, and answers to the extraction points for each of the included papers were obtained by two different authors, resolving differences by discussion between the two involved authors; disagreements were solved by discussion within the entire author consortium. While extracting, we distinguished between explanation methods and model interpretability according to the nomenclature above introduced, even though some papers did not agree with the latter.
A flowchart summarizing the number of papers found, excluded, and included in this study, along the lines of [40], is displayed in Fig. 2.
Figure 2: Flow diagram indicating the number of papers identified, excluded in the different phases, and included in the review.
IV Interplay between FL and XAI
This section explores studies that have either integrated explainability into FL methods or analyzed how one technology influences the other. Studies that jointly deal with FL an XAI but do not necessarily report an influence are discussed in the Appendix. Here, we first discuss instances where FL has impacted post-hoc explanations (Sec. IV-A). Next, we examine works where explanation methods have influenced FL training as a design step in the FL process (Sec. IV-B). Lastly, we cover studies where FL has altered model interpretability, or the implementation of interpretable models has influenced FL training (Sec. IV-C).
IV-A FL impacting XAI
Among all papers reviewed, only [41] quantitatively analyzes the impact of FL on explanations. Additionally, [42] highlights privacy concerns associated with an explanation method used on FL models.
An FL procedure is developed in [41] for taxi travel-time prediction, based on time-series and geographical data. The authors develop a federated feature attribution aggregation method and test the similarity of FL-calculated explanations compared to those calculated centrally. Several XAI techniques are tested, specifically DL-specific attribution explanation methods such as DeepLIFT [43], Saliency [44], Input X Gradient [45], Guided Backpropagation [46], Deconvolution [47], and Layer-wise Relevance Propagation [48]. It is reported that, upon training via FL, explanations across different local models lead to different variable importance rankings. However, no significant difference was found in federated feature attributions in FL compared to those in centrally trained models.
An FL model is trained in [42] to predict the latency in the creation of a network slice in a mobile communications scenario. Each slice manager provides data regarding CPU/RAM capacity and usage, serving as an FL node. The models are evaluated both locally and globally using SHAP [20], LIME [49], Partial Dependence Plots (PDP) [50, Ch. 8.1], and RuleFit [51]. For the PDP technique, the FL clients only share plots that display the percentage of feature impact on the target label to preserve data privacy. The influence is analyzed qualitatively with a focus on privacy concerns. Overall, it is shown that PDP explanations raise privacy concerns since they are executed on the client side.
IV-B Explanation method impacts FL training
Among the papers that reported that the use of an explanation method impacts the FL training, a salient subset uses XAI to defend the FL process from the negative effects of defaulting nodes [33], malicious behavior from certain FL nodes [52, 53], and instances of the GAN attack in FL [54]. Other positive effects reported include improved accuracy in [55, 36, 56] and enhanced learning efficiency [57]. It must be noted that such benefits stem from incorporating XAI as a component of the FL algorithm design.
With the aim of detecting malicious attacks on FL operations, [53] uses a random forest (RF) to identify features causing incorrect predictions. Each participant trains both a DL and an RF model on their training data. For samples that are misclassified, it calculates the feature importance of each feature regarding incorrect classification, from all decision trees and using LIME [49]. The change in a feature’s importance over time is used to assign the contribution of each feature in misclassifying the data. Such an importance value provides insights into the level of influence of each log key in a sequence during an attack, thereby indicating which features should be most protected.
A novel FL protocol was proposed in [36] where a subset of all centers available online were selected to participate in each FL round, using Shapley values (computed using SHAP) as a heuristic to estimate FL contributions from each client. More specifically, at each round, the difference between local (feature-wise) and global aggregated Shapley values at each node is used to select participating parties in an FL process. The paper reports that the proposed method improves both the efficiency and accuracy compared to the FedAvg protocol [10], which does not consider client contributions.
The idea in [52] is to use explanation methods (specifically Backpropagation, GuidedBP, DeepLift, GradCAM [39], and Integrated Gradients) to detect whether each participant is using malicious data to enact a backdoor attack. To this end, so-called detection filters are developed. These consist of a classifier and an explanation method that respectively identify a likely backdoor attack and triggering features in the input data. The effectiveness of various explanation methods with different classifiers is tested, strengthening the FL process against backdoor attacks. Upon testing the detection accuracy in the presence of variable proportions of backdoor attacks, the proposed methodology proves the usefulness of the different explanation methods for backdoor attack prevention.
The technique proposed in [56] uses Shapley values and the Lipschitz constant to generate both local (feature-wise) and global explanations. Local Shapley values are compared with global Shapley values to refine the training of the local model, ensuring that only necessary characteristics are retrained, which allows for the personalization of the FL model for each user so that only the necessary characteristics of the model are retrained. A rigorous methodology is applied, resulting in increased accuracy, quantified explanations, and reduced dependence on input shift.
In an image classification task, Shapley values calculated via SHAP are used in [54] to identify the most important pixels for each local FL model and mask the pixels with the highest SHAP scores. The resulting dataset is then used for training the FL model. This method is proposed to protect the FL setup from adversarial attacks, specifically poisoning GAN attacks. By masking the majority of influential pixels in the input images, the difficulty of the GAN attack on FL is increased (although the exact influence of the method is not quantitatively assessed).
The algorithm proposed in [55] is designed to explain the output of a time-series classifier. It extracts input subsequences that highly activate a time-domain convolutional neural network, facilitating their visualization. A graph capturing temporal dependencies is computed at each learning node. The central server aggregates these graphs into a global temporal evolution graph. By applying this method, an improved classification accuracy is claimed, compared to other FL algorithms such as FedAvg, FedRep, and FetchSGD.
The method proposed in [33], termed Node Liability in Federated Learning (NL-FL), traces back ML decisions to training data sources in distributed settings. This method allows for the identification of misbehaving (defaulting) nodes that can be excluded from the training process. The influence of each node is quantified by measuring the classification accuracy after removing misbehaving nodes. The proposed method results in an improved prediction accuracy.
Adaptive sparse deep networks are implemented in [57], where parameters are shared via a multi-level federated network. At each round, weights are shared at the “top” and “second” sharing levels of the FL architecture, depending on the relevance values of the network calculated through Layerwise Relevance Propagation (LRP) [58]. This approach provides good diagnostic results even when the FL dataset is under a non-independent identical distribution (NOIID).
IV-C FL impacting model interpretability and vice versa
This subsection examines the relationship between Federated Learning (FL) and model interpretability. The featured studies, [59] and [60], discuss both the benefits and challenges of interpretability in distributed learning environments. These examples illustrate the trade-offs between model complexity and transparency that are inherent in FL setups.
An aggregation of client-based attention weights is investigated in [59] for a threat-detection task in a cloud scenario. Using system logs as input data, each client predicts cyberattacks and computes local attention weights, which are claimed to enhance interpretability. The central server subsequently aggregates these attention weights to build a saliency map that provides insights into the impact of the different log keys on the threat prediction. The influence of FL on model interpretability is assessed by comparing insights and interpretations from attention mechanisms for individual cyberattacks per client, individual attacks per aggregated models, and aggregated attacks per aggregated model. The aggregated insights proved to be more general, but at the expense of reduced interpretability.
In [60], a fuzzy rule-based system (FRBS) [61] is trained in federation via a one-shot communication scheme. Each data silo computes their own FRBS, and the individual models are merged by the central server. The proposed FRBS uses a maximum-matching inference rule so the inferred regression function is piecewise linear, which is inherently explainable. It is observed in this work that FL and interpretability impact each other. The authors report that the average number of model rules is higher when trained in the federated setting (vs. centralized). Following the intuitive idea that higher complexity implies lower interpretability, this means that training via FL implies a lower interpretability in this specific study. On the other hand, the inherent interpretability of the models used in this study’s approach (Takagi-Sugeno-Kang FRBS [62]) allows the central server to identify conflicts between rules inferred by data centers. This approach facilitates the reconciliation of such discrepancies, thereby enhancing the learning process. This reconciliation not only improves model performance but also results in a more interpretable model compared to the ensemble of local models.
V Results
In this section, the information extracted from the papers included in the review is presented in a graphical form, to provide an overview of the different fields of application, explainability techniques, FL setups, and data types and modalities.
Figure 3: Field of application.
Fig. 3 shows the proportion of papers across the different fields of application, where medicine and life sciences are predominant (40.5%), followed by finance (16.2%), cybersecurity (13.5%), telecommunication (8.1%) and optimization in other engineering fields (8.1%) such as electricity, mechanics, and transportation. A small proportion (13.5%) of the analyzed papers focused on the theoretical side of the ML techniques without mentioning any specific application.
Figure 4: Explainability techniques. Shades of green represent interpretable models; shades of blue represent explanation methods.
Technique Description Manifestations in sample
Feature relevance Calculate relevance scores for model variables. e.g., SHAP [36, 41, 56],
Custom builds [55, 63]
Local explanations Estimate whole model through less complex subsystems. e.g., GradCAM [64, 52, 65],
heatmaps [66]
Simplification Facilitate model while maintaining performance. e.g., simpler models [67, 60]
Text explanations Generate symbols that explain the results of the model. -
Visual explanations Visualize the inference process of the model. -
Explanations by example Provide representative examples that allow insight into the model. -
Algorithmic transparency Enable users to follow and understand the processes by the model. e.g., Linear regression [68],
Decision trees [69, 70]
Decomposability Explain each model part separately for full comprehension. e.g., Inference splits [71]
Simulatability The inference of models could be simulated by a human. e.g., Rule-based systems [72]
TABLE II: Brief descriptions of the explainability techniques used in this review as understood in [18].
The proportion of interpretability and explanation methods is shown in Fig. 4 and described in Tab. II. The most frequently used type of interpretability of FL models is algorithmic transparency, with some use of decomposability and simulatability [18]. While algorithmic transparency allows users to algorithmically trace a model’s processes, decomposability describes models of which fractions can be understood by humans. Simulatability refers to models that are sufficiently interpretable for a human to understand, or “simulate”, as a whole [18]. Among the explanation methods applied with FL models, feature relevance is predominant, followed by local explanations, and a smaller amount of model simplification-based methods [18]. Feature relevance methods quantify the impact of the model’s input features, local explanations explain specific model predictions, and simplification refers to rebuilding the entire model for easier explainability.
Agreement yes no - SUM
Definition
yes 10 7 0 17
no 7 4 9 20
SUM 17 11 9 37
TABLE III: Nomenclature and definitions.
Table III shows the amount of reviewed papers where a definition of XAI is provided and for which of these there is an agreement between their notion of XAI and the one expressed in Sec. I. Out of the 17 papers where a definition is given, 10 agree with ours and 7 do not. When a definition is not given, in 9 cases it was not possible to infer what notion of XAI is used, 7 managed the term in accordance to our definition and 4 in disagreement.
Figure 5: Amount of data used for each FL setup.
Fig. 5 shows the distribution of papers using different amounts of data for the experiments, while the colors indicate whether the FL setup was real, simulated or not specified. No papers in this survey reported real FL setups using more than 1,000,000 data points. Out of the 7 FL setups using between 100,000 and 1,000,000 data points, 3 did so in a real FL setup [73, 70, 74].
Figure 6: Number of data centers in the FL networks for each setup.
Fig. 6 shows the distribution of papers using different amounts of FL centers, while the colors indicate whether the FL setup was real, simulated, or not specified. The highest number of nodes (1,000) was simulated by [52] using standard benchmark data (MNIST, CIFAR10), followed by [9] with 100 nodes, using text datasets for language modeling; [53] with 100 nodes, using physical activity and census data from the UCI ML repository; and [75] with 66 nodes, using data from a cyberattack classification task. A minority of 9 out of 37 studies used real setups, where [76] used 12 nodes and all other papers implementing a real FL setup used 10 or less nodes. 11 papers used between 3 and 10 nodes, 2 of which using a real setup; and 9 papers used 3 or fewer centers, 5 of which using a real setup.
Figure 7: FL libraries.
Only 27.1% of the papers included in this survey used established FL libraries, whereas 37,8% developed their own libraries, and 35.1% did not specify how the network was implemented. Fig. 7 shows the FL libraries used by the analyzed papers.
Figure 8: Type of FL used.
Fig. 8 illustrates the distribution of FL types used in the reviewed papers. The most prominent is horizontal FL, with 30 studies using this type exclusively. The vertical FL type is used far less, appearing in only four studies, and transfer FL is the least used, with only one study focusing on it. The overlaps in the diagram show that few studies used multiple of these methods. Only one study uses horizontal and vertical FL, another uses horizontal and transfer FL, another uses vertical and transfer FL, and one uses all three types.
Figure 9: type of data used in experiments.
Figure 10: Modality of data.
Regarding the type and modality of data, Fig. 9 shows that a majority (30%) of works experimented with real clinical data followed by a 17.5% using real finance data. 15% used widespread datasets typically oriented to model benchmarking, and the remaining works used a diversity of real data types. According to Fig. 10, the predominant modality is tabular data (52.8%) followed by images (19.4%) and time series (22.2%).
Figure 11: Number of papers that reported influence of FL on XAI compared to the amount of data used.
Fig. 11 illustrates the number of papers that reported any influence between FL and XAI in relation to the amount of data used in the experiments. None of the papers using less than 1,000 data points reported any influence, whereas 9 papers working with more than 1,000 data points reported an influence.
VI Discussion
In our sample of papers jointly investigating FL and XAI,we found mostly post-hoc explanation methods. The application of explanation methods on FL systems is predominantly based on feature importance followed by local explanations, reflecting the prevalence of such methods in the literature. None of the reviewed works have proposed an ad-hoc FL method tailored to the ulterior application of explanation methods, highlighting the need for future approaches in that domain. Explanation methods designed for federated trained ML models were limited to feature aggregation [41], control of information sharing [42], and counterfactual explanations in VFL [77]. The works dealing with FL of interpretable models focused mostly on algorithmic transparency, and in the case of [60] the proposed method learns a set of interpretable rules that reflect the structure of the FL network. Interpretable models trained with FL methods were limited to fuzzy rule-based systems [60], time-series classifiers [71], SVM [78], Cox proportional hazards [79], sparse Bayesian models [80], and decision trees [69]. Fewer works deal with decomposability or simulatability, probably due to an increased difficulty in imposing those properties in an FL system.
Regarding novel methods to train interpretable ML models via FL, one very promising approach allows building federated classification models without relying on gradient descent-based methods [67]. This differs from most methods in the current literature, which focus on differentiable models. The approach of [67] is based on adapting previously existing algorithms in the boosting family to FL. Other approaches proposed an optimization method to solve the sparse SVM problem in FL [78], a novel technique called Vanishing Boosted Weights [81], or an FL network based on gradient-boosting decision trees [75].
A preference for HFL is observed in the studies surveyed, while combinations of FL types and the use of TFL are relatively rare. This is in accordance with the observed higher prevalence of HFL in the literature compared to VFL [82, 83]. Moreover, most papers apply cross-silo FL with a small set of data centers. Whether this is due to data access or this is an accurate representation of FL networks remains unclear. In general, experiments on the effect of FL on interpretability were very sparse. [60] observed an influence of FL on model interpretability in the FRBS trained using their own method, namely a larger number of model rules. Similarly, in [59], by interpreting the attention weights produced by transformer models based on a transformer architecture [84], different insights for each client were observed.
The fact that a majority of the reviewed papers focus on healthcare, finance, and engineering applications such as networking, highlights a high degree of interdisciplinarity and suggests cross-fertilization among fields. Conversely, this contrasts with the relative lack of studies addressing high-stakes applications in social networks, language models, and supply-chain management, where user privacy and transparent decision-making are also crucial. The need to define explicability and privacy requirements usually originates from the end users’ perspective, where data interoperability and reasoning behind automated decisions are key. The implementation of enabling technologies originates research towards using XAI for more technical goals such as improving model accuracy, assessing the training process, and prevent malicious behavior, as observed in Sec. IV-B.
In several reviewed cases, aggregating interpretability metrics from different nodes led to more general but less interpretable global insights. This fact supports the idea that FL promotes generalizability at the cost of interpretable insights from the local nodes of the FL framework. More specifically, aggregating feature relevance or attention weights across nodes can yield more generalized global insights but often sacrifices some degree of localized interpretability.
This was reported in the case of an attention mechanism [59] where each client trained locally on system log data and generated local attention weights. The global saliency map resulting from aggregating the local attention weights from the different nodes, provided a global saliency map where the contribution of individual client models is diluted, possibly leading to unique local patterns getting lost during global aggregation. The FRBS applied by [60] in a federated context involves rule generation at each data silo and merging them at the central server. Experiments observed an increased number of rules in the global model compared to local ones, indicating a growth in model complexity upon aggregating the local models, yielding a more expressive global model that is less interpretable than individual local models.
This result suggests that while FL can improve overall model performance and integrate information from diverse data sources, outputs from the global FL model could become more difficult to interpret than those from local models. This may negatively affect applications where the transparency of the local models is beneficial. One example is medical prognosis, where each local model is associated with a clinical site or hospital, and the distributions of disease features vary across hospitals. This issue could benefit from new aggregation strategies or federated explanation techniques that retain node-specific insights without compromising global model integrity.
Despite the prevalence of SHAP among the included studies, we did not find any mention of a potential federated implementation of SHAP when the supporting dataset is distributed among several data centers. Such a contribution would help exploit the representativeness of supporting data from diverse centers, potentially helping generate more accurate Shapley values. Among the studies using small amounts of data, very few reported an influence of FL on XAI, suggesting that an insufficient amount of data makes these observations challenging. It should be noted that popular feature attribution methods such as SHAP require a set of “supporting” data to do the approximate computations. Ideally, such data should not be part of the training dataset, so an insufficient number of data points probably does not allow generation of explanations robust enough to draw conclusions.
Most studies were done with simulated FL networks, and the ones done with a real FL setup used datasets with less than one million samples. With ever-increasing models trained on billions of data points, the practical implications of big data remain unexplored in the context of FL and XAI. Similarly, a minority of the reviewed works use established FL libraries. The remaining papers either develop their own libraries or lack descriptions of how the network was implemented. The absence of standardized reporting for used libraries can lead to misleading observations if there are flaws in the libraries that are not identified during the analysis. A lack of transparency can hinder the development of standardized solutions, as inconsistencies in the reporting and usage of libraries complicate replicability. We also observed that many articles could improve their way of reporting data characteristics since strict adherence to reporting guidelines such as TRIPOD+AI [85] or MINIMAR [86] would ensure reproducibility and transparency. This can help standardize FL implementations, potentially accelerating the adoption of best practices in FL and XAI.
In the same line, not providing a definition of explainability [cf. Table III] can be a hurdle for reproducibility and further analysis. Works often disagree on what should be referred to as interpretability or an explanation method, which we aim to align in this study. Clarity and consensus in nomenclature are of utmost importance because they foster consistency in the evaluation of the understandability of ML results, improve communication among researchers and policymakers, and enable the development and reproducibility of new XAI methods.
The scarcity of studies explicitly addressing and quantifying the influence of FL on explanations does not give grounds for concluding that there is not such an influence, revealing an interesting research gap. Principled experimental work aimed at objectively evaluating the impact of FL training on ML model structure is also expected to shed light on its consequences on model explanations and interpretability.
A limitation of this study and suggestion for future work is that the used FL algorithm (e.g. FedAvg) was not extracted, which could be a relevant factor in analyzing the impact on explanations. As stated above, adherence to information reporting is important, and future reviews on the field should look closer at this aspect. Further, we conducted our search in 2023. In the ever-changing domain of ML model development, further developments could have been achieved in the meantime.
VII Conclusion
This review highlights the interplay between FL and XAI, both methodologically and experimentally. We have identified a research gap in which few studies quantify the impact of FL on explanations. Moreover, the impact of FL on model interpretability and explanations remains unclear, revealing the need for studies that quantify such impact. There is a need for rigorous experimental and analytical research to assess how FL training influences the structure of the model and its implications for explainability and interpretability. Additional research should also provide guidance for practitioners on the responsible implementation of FL with XAI, particularly in critical fields such as healthcare, finance, and engineering. This guidance will help ensure responsible deployment of AI systems that balance model performance with due transparency.
Many papers do not use consistent terminology for explainability, which complicated the analysis. It is important for future research to clearly define terms and use standardized nomenclature. Another important finding is that a minority of the studies specify which FL libraries were used. Future publications should include details on the libraries employed or provide source code for custom implementations. Furthermore, the lack of adequate reporting of data characteristics calls for strict adherence to reporting guidelines. These are important improvements towards ensuring reproducibility, transparency, and auditability, which are crucial ethical aspects in sectors such as medicine.
In conclusion, this review underscores the need for more structured and transparent research practices in the intersection of FL and XAI. Establishing clear definitions and consistent methodologies will be key in advancing the field and addressing the identified gaps. As demand for FL and XAI continues to grow, particularly in high-stakes environments, the importance of rigorous, transparent, and reproducible research cannot be overstated.
Appendix A Extended List of Reviewed Papers
In this section, we describe how each selected paper deals with FL and XAI in combination. The first part of the section discusses works where FL is combined with explanation methods, both methodological (Sec. A-A) and applied (Sec. A-B). The second part of the section discusses papers combining FL and interpretable models, both methodological (Sec. A-C) and applied (Sec. A-D).
A-A Combining FL and explanation methods: methodological contributions
[36] proposes a novel FL protocol where a subset of all centers available online are selected to participate in each FL round by calculating the difference between the aggregated global and local feature contribution (using SHAP [20]). A more efficient and improved performance is shown when using this selection.
[87] proposes the first counterfactual explanation method for VFL. They show the validity by retraining VFL models on banking data while leaving out a varying number of features with respect to their counterfactual importance rate and comparing against random selection of variables.
[52] uses explanation methods such as GradCAM [39] to detect whether each participant is using malicious data to enact a backdoor attack by developing so-called detection filters. These consist of a classifier and an explanation method that identify a likely backdoor attack and triggering features in the input data, respectively. The effectiveness of various explanation methods with different classifiers is tested, strengthening the FL process against backdoor attacks.
[54] tackles an image classification task by calculating Shapley values by SHAP and using them to find the most important pixels for every local FL model and mask the pixels with highest SHAP score. The resulting dataset is used for training the FL model. This method is proposed to protect the FL setup from adversarial attacks, specifically poisoning GAN attacks.
[53] uses random forest (RF) algorithms to detect features causing wrong predictions, with emphasis on detecting malicious attacks on the FL operation. Each participant trains a DL and RF model on its training data. For the samples that are wrongly classified, it calculates the average feature importance of all decision trees with LIME [49] that will result in the same wrong classification. It uses the change in this feature’s importance over time to predict the contribution of each feature in wrongly classifying the data.
[55] proposes and designs an algorithm for explaining the output of a time-series classifier. It extracts and visualizes the input subsequences that highly activate a convolutional neural network. A graph capturing temporal dependencies is computed at each learning node. The central server aggregates the obtained graphs into a global temporal evolution graph.
[88] introduces PrADA, a privacy-preserving federated adversarial domain adaptation technique addressing cross-silo federated domain adaptation issues. PrADA mitigates sample and feature scarcity by employing VFL with a feature-rich party and implementing adversarial domain adaptation from a sample-abundant source. For interpretability, features are segregated into semantically meaningful groups for fine-grained adaptation based on Shapley values computed using SHAP.
[33] proposes a method, namely node liability in federated learning (NL-FL), to trace back ML decisions to training data sources in distributed settings. The method allows for the identification of misbehaving nodes that can be excluded from the training process, resulting in improved prediction results.
[57] implements interpretable adaptive sparse deep networks which exchange NN parameters by means of a multi-level federated network. Whether those weights are shared at the “top sharing level” of the FL architecture depends on the relevance values of the network calculated through layerwise relevance propagation (LRP). The approach provides good diagnostic results even when the FL dataset is under a non-independent identical distribution (NOIID).
A-B Combining FL and explanation methods: applied contributions
[64] trains FL models to segment lung X-ray images and detect signs of pneumonia. Grad-CAM is used to highlight parts of the images that contribute to a detection. It is concluded that a model trained on segmented images has less accuracy, but the pixels highlighted by Grad-CAM focus more on the lung area. It is also reported that training the model in the FL manner helps maintain generalizability and avoid overfitting. A fixed number of FL rounds and greater number of local iterations result into more accuracy.
[63] aims to predict residential load using a recurrent neural network (RNN). To explain the importance of features, the authors propose a novel automatic relevance determination (ARD) method. An iterative federated clustering algorithm (IFCA) is used, which keeps several central models while clustering the input data sequences, and each model is updated using data in its associated cluster. No interaction between ARD and IFCA is reported.
[41] develops an FL procedure for taxi travel-time prediction based on time-series and geographical data. Authors develop a federated feature attribution aggregation method and test how similar the FL calculated explanations are compared to central calculation. Many XAI techniques are tested, all result in similarly low differences.
[56] compares the use of Shapley values and Lipschitz constant for generating both local and global explanations, and uses this information to update the model. It allows for the personalization of the FL model for each user, so that only the necessary characteristics of the model are retrained, based on the respective needs and the events that it is called to respond.
[89] uses Shapley values computed via SHAP to explain outputs of an FL model trained on edge devices to its operators. It takes the model and the test data as inputs to construct a local linear regression explanation model. Subsequently, the explanatory model computes the Shapley values of classified anomalies and displays them visually. As feature values are measured by sensors, the explanations help operators determine the sensors likely causing an abnormality and make a faster detection response.
[42] proposes to train an FL model to predict the latency in the creation of a network slice. Each slice manager provides data regarding CPU/RAM capacity and usage and serve as FL nodes. The models are evaluated on a local and global level using SHAP, LIME, partial dependent plot (PDP) [50, Ch. 8.1] and RuleFit [51]. Overall, they show that PDP explanations raise privacy concerns since they are run on the client side.
[90] discusses the use of FL and transfer learning to enhance AI-based lung segmentation. The study achieved good segmentation accuracy using local system data and pre-trained weights from U-net models. The approach utilized a reduced number of nodes with varying dataset sizes and incorporated a model-agnostic explanation method (activation map) to clarify the results.
[73] trains a language model that takes free text from electronic medical records and classifies a patient’s disease into one of the International Classification of Diseases (ICD-10) codes. They showed explanations for the predictions by highlighting input words via a label attention architecture. FL is realized via Flower [91] to integrate training data from 3 different sites while keeping data privacy.
[92] introduces AnoFed, a framework integrating transformer-based Autoencoders (AEs) and Variational Autoencoders (VAEs) with Support Vector Data Description (SVDD) in a federated environment, specifically for ECG anomaly detection, optimizing computational efficiency. A combined design of the VAE/AE and SVDD incorporates kernel density estimation for adaptive anomaly detection. Moreover, it includes an module which explains the anomaly detection output by identifying key segments of the ECG signal that show the maximum reconstruction loss.
[74] presents an ECG-based arrhythmia classification framework which trains convolutional DNNs via FL, and includes an XAI module computing activation mappings in the ECG signal by means of GradCAM. The framework addresses data availability, privacy, and interpretability challenges.
[66] proposes a framework for FL in connected medical devices with blockchain integration for safe model parameter exchange. That framework is showcased by a hybrid implementation of actual hardware nodes and simulated distribution of publicly available datasets in many use cases. Explanations are shown in two cases, however the results and impact of FL are not discussed.
A-C Papers combining FL and interpretable models: methodological contributions
[67] proposes an FL algorithm that builds federated classification models without relying on gradient descent-based methods. Therefore, the class of algorithms that can be learned via FL is not restricted to models whose output is differentiable with respect to the model parameters. Based on AdaBoost [93], it effectively combines gradient-free classifiers which may be learned independently by the FL clients.
[60] trains a fuzzy rule-based system (FRBS) [61] in federation via a one-shot communication scheme where each data silo computes their own FRBS and the individual models are merged by the central server. The proposed FRBS uses a maximum-matching inference rule so the inferred regression function is piecewise linear, which is inherently explainable.
[71] proposes an FL framework for time-series classification , using interpretable, human-understandable time-series features, namely shapelet features, interval features, and dictionary features. The paper claims to guarantee interpretability for the learning-initiating party by ensuring that it can access the aforementioned features without data leakage. To ensure security, the solution incorporates secure feature extraction protocols, secure model training protocols, additive secret sharing schemes, and secure computation protocols.
[72] highlights the importance of using information granules for better interpretability, focusing on unsupervised federated learning and enhancing rule-based models through granule decomposition and linguistic approximation.
[70] proposes an Efficient and Interpretable Inference Framework for Decision Tree Ensembles in FL (Fed-EINI), based on streamlined multi-party communication. The paper highlights the challenge of current privacy-preserving ML frameworks compromising model interpretability to prevent data breaches. The proposed solution enhances interpretability by disclosing feature meanings while maintaining privacy.
[76] presents an interpretable data interoperability method for FL called iFedAvg to address the low interoperability due to client data inconsistencies, among other challenges. The iFedAvg method uses personalized layers to adjust for local data shifts, like age differences, directly within input features, which maintains privacy while allowing for direct interpretability. The difference in values of private weight and bias of input layer of each participant captures the inherent shift in data. It was tested on public benchmarks and on a large, real-world Ebola dataset.
[94] introduces a group personalization strategy in FL to address client drift in settings with distinct client partitions. Authors fine-tuned a global FL model with another FL process for each homogeneous client group and then adapted it per client. The method is tested on real-world language modeling datasets and aligns with Bayesian hierarchical modeling principles.
[69] introduces an interpretable FL system for collaborative data analysis across distributed networks using interpretable models such as decision trees, sharing intermediate representations of the data rather than models. The result is an interpretable model that performs better than individual analyses and nearly as well as centralized methods.
[65] introduces adaptive differential privacy (ADP) in FL to balance privacy and model interpretability, assessed by inspecting Grad-CAM heatmaps. ADP selectively injects noise into client model gradients, mitigating gradient leakage attacks while preserving interpretability. Through theoretical and experimental analyses on IID and NOIID data, it overcomes the limitations of traditional differential privacy, demonstrating a harmonious blend of data privacy safeguards and interpretability.
[78] proposes a framework for solving sparse support vector machine (SVM) classification in a distributed fashion. Its performance demonstrated on an electronic health record (EHR) dataset. The proposed algorithm has improved convergence rate compared to several alternatives. Interpretability is assessed by achieving a classifier with fewer features considered as highly important for predictions.
A-D Papers combining FL and interpretable models: applied contributions
[75] implements a FL network based on Gradient Boosting Decision Trees (GBDT). These GBDT are considered transparent models and therefore, their FL network is considered transparent as well. They apply their algorithm on network intrusion data sets and claim that model is interpretable for human experts.
[68] introduces a method in financial risk management for credit scoring and rating with big-data capabilities. Using a VFL framework, it allows multiple agencies to collaboratively train an optimized scorecard model. Performance is showcased on two finance datasets.
[81] proposes a novel technique called Vanishing Boosted Weights for fine-tuning models trained by a GBDT algorithm, along with an FL version of this approach. Based on stumps, the model remains interpretable due to their limited number. Iterative adjustment of each stump’s output value occurs multiple times by incorporating a vanishing sequence of values.
[59] investigates client-based attention weight aggregation in a threat-detection task within a cloud scenario, using system logs as input data. Each client predicts local attentions, which are claimed to enhance interpretability, and the central server subsequently aggregates the attention weights to build a saliency map that provides insights on the impact of the different log keys on the threat prediction.
[79] proposes an adaptation for FL of the Cox Proportional Hazards regression model with LASSO regularization. Such an estimator is used as a feature selector in the context of survival analysis and personalized medicine. The inclusion of LASSO regularization enacts a feature selection, contributing to the model’s interpretability.
[80] proposes methods for training sparse Bayesian models in federation, allowing pooling from multiple data sources without privacy issues and offering principled uncertainty quantification. The methods are based on Markov Chain Monte Carlo (MCMC) updating steps, where the order of updating steps can be interchanged so the communication between local servers and the global server can be reduced by running multiple local steps per global aggregation.
[95] explores FL of interpretable models in 5G and 6G systems, focusing on automated vehicle networking. The approach addresses gaps in existing AI-based solutions for wireless networks, particularly in vehicle-to-everything (V2X) environments. The methodology offers decentralized, efficient intelligence, enhancing operational trustworthiness and data management.